OpenStudy (gavin39):

What is the simplified form of the quantity 1 over y plus 1 over x all over 1 over y minus 1 over x ? y plus x, over y minus x y minus x, over y plus x x minus y, over x plus y x plus y, over x minus y

OpenStudy (gavin39):

Need a quick anwser

OpenStudy (gavin39):

@Mertsj

OpenStudy (mertsj):

\[\frac{\frac{1}{y}+\frac{1}{x}}{\frac{1}{y}-\frac{1}{x}}=\frac{x+y}{x-y}\]